Question 10 When 1.00 g of thiamine hydrochloride (also calledvitamin B1 hydrochloride) is dissolved in water, then diluted to 10.00 mL, the pH of the resulting solution is 4.50. The formula weight of thiamine hydrochloride is 337.3 g/mol. Calculate K Not answered Marked out of 1.00 Answer: Flag questionCan you please help me solve this, I'm not sure how to solve this one. I will rate well, thank you!
Answer #1

molarity = W*1000/G.M.Wt* volume of solution in ml

               = 1*1000/337.3*10   = 0.296M

PH   = 1/2[Pka-logc]

4.5    = 1/2[PKa-log0.296]

9       = [PKa-log0.296]

9      = Pka +0.53

PKa   = 9-0.53 = 8.47

PKa   = 8.47

-logKa = 8.47

    Ka   = 10^-8.47   = 3.4*10^-9